# Discovering modular solutions that generalize compositionally

### v1

@@ Abstract
Many complex tasks and environments can be decomposed into simpler, independent parts. Discovering such underlying compositional structure has the potential to expedite adaptation and enable compositional generalization. Despite progress, our most powerful systems struggle to compose flexibly. While most of these systems are monolithic, modularity promises to allow capturing the compositional nature of many tasks. However, it is unclear under which circumstances modular systems discover this hidden compositional structure. To shed light on this question, we study a teacher-student setting with a modular teacher where we have full control over the composition of ground truth modules. This allows us to relate the problem of compositional generalization to that of identification of the underlying modules. We show theoretically that identification up to linear transformation purely from demonstrations is possible in hypernetworks without having to learn an exponential number of module combinations. While our theory assumes the infinite data limit, in an extensive empirical study we demonstrate how meta-learning from finite data can discover modular solutions that generalize compositionally in modular but not monolithic architectures. We further show that our insights translate outside the teacher-student setting and demonstrate that in tasks with compositional preferences and tasks with compositional goals hypernetworks can discover modular policies that compositionally generalize.  [^1] Code available at https://github.com/smonsays/modular-hyperteacher
@@ 1 Introduction
While monolithic architectures, in which the whole network is engaged for all tasks, are powerful, they lack a built-in mechanism to decompose learned knowledge into modules. This puts them at a potential disadvantage as they might have to learn each of the exponentially many combinations individually even if their constituent modules were all previously encountered. For instance, it is unclear to what extent monolithic large language models can compositionally generalize beyond the training distribution (Srivastava et al., 2023; Press et al., 2023; Dziri et al., 2023).
This raises the question whether a modular architecture built from independent modules that can be combined flexibly can more easily extract modular structure from an environment with underlying compositional structure. Typically modular architectures are expressive enough to collapse to the monolithic solution of modeling all task combinations encountered during training and often do so in practice (Kirsch et al., 2018; Shazeer et al., 2017; Bengio et al., 2016; Rosenbaum et al., 2018; Mittal et al., 2022). Identifying the correct modules is further complicated by the fact that the learning system only ever encounters data generated from mixtures of modules without access to the ground truth decomposition.
Here we aim to illuminate the question what properties of the data generating process allow to discover modular solutions enabling compositional generalization. We do so within the framework of multi-task learning where a learner encounters input/output pairs on a set of tasks composed of few underlying modules. As modular architecture, we specifically consider hypernetworks (Ha et al., 2017), a general form of multiplicative interaction (Jayakumar et al., 2020) which allow to express compositions in weight space used by many multi-task models (Kumar & Daume III, 2012; Ruvolo & Eaton, 2013; Perez et al., 2018). Our main contributions are as follows:
@@
  • In an extended teacher-student setting, we demonstrate empirically how meta-learning from finite data of sparse module combinations can discover modular solutions that generalize compositionally in modular but not monolithic architectures.
  • We validate that our results translate to more realistic settings of learning action-value functions for compositional preferences and learning of behavior policies for compositional goals.
Figure 1: A Diagram contrasting the hypernetwork as a modular architecture to a monolithic architecture. B Toy example illustrating the modular teacher-student setting where tasks are drawn from parameter module compositions of a teacher hypernetwork.
@@ 2 Methods
To endow our data generating distributions with compositional structure, for each task we recombine sparse subsets of up to $K$ components from a pool of $M$ modules. Each task can be fully described by a latent code ${\bm{z}}\in\mathbb{R}^{m}$ that specifies the combination of module parameters $(\Theta^{(m)})_{1\leq m\leq M}$ which parameterizes a task-shared function $g$ with parameters $\bm{\omega}({\bm{z}})=\sum^{M}_{m=1}z^{(m)}\Theta^{(m)}$. Being exposed only to demonstrations of such tasks with their underlying modular structure hidden, the naive strategy for a learner would be to learn every combination separately inside a single monolithic network. As the number of modules grows, the quantity of possible combinations increases exponentially, requiring more and more capacity. In contrast, identifying underlying compositional structure not only affords a more compact representation, it also potentially allows the learner to generalize to unseen compositions.
@@
For a learning system to be able to leverage the compositional structure of a family of tasks, its model architecture needs to be able to efficiently represent the corresponding class of compositions. Here we contrast modular and monolithic architectures (see Figure 1A) and investigate under which circumstances they admit functionally modular solutions that reflect the compositional structure of the data generating distribution. To test our models for compositional generalization, we hold-out a number of module combinations when we sample tasks for training. At evaluation time we can then test the ability of all models to solve the held-out out-of-distribution (OOD) tasks.
As our modular architecture of choice we consider hypernetworks (Ha et al., 2017) that allow to express a wide class of multiplicative interactions (Jayakumar et al., 2020). A hypernetwork $h$ is a neural network that generates the parameters $\bm{\omega}$ for a target neural network $g({\bm{x}};\bm{\omega})$ which itself processes the task data. The hypernetwork takes as input a task-specific embedding vector $\bm{\phi}$ and is parameterized by the task-shared parameters $\bm{\theta}$, i.e. $\bm{\omega}=h(\bm{\phi};\bm{\theta})$. We can think of $h$ as a function that maps $\bm{\phi}$ to a selection of flattened module parameters $(\Theta^{(m)})_{1\leq m\leq M}$. In the case of a linear hypernetwork, $h$ is simply a linear function of $\bm{\phi}$ resulting in a linear combination of the module parameters. In the case of nonlinear hypernetworks, $h$ is a separate neural network whose penultimate layer can be thought of as selecting among the module parameters.
As an example for a popular class of monolithic architectures we consider the commonly used motif of combining a large and expressive feature encoder shared across tasks combined with a task-specific decoder (e.g. Lee et al., 2019; Bertinetto et al., 2019; Raghu et al., 2020). Here we focus on ANIL (Raghu et al., 2020) which uses a linear decoder, i.e. $f({\bm{x}},\bm{\phi},\bm{\theta})=\bm{\phi}^{\top}g({\bm{x}};\bm{\theta})$ where $\bm{\phi}$ is a task-specific matrix. In addition, we consider the widespread MAML architecture (Finn et al., 2017) with $f({\bm{x}},\bm{\phi},\bm{\theta})=g({\bm{x}};\bm{\phi}(\bm{\theta}))$ where the task-shared parameters $\bm{\theta}$ are used to initialize the task-specific parameters $\bm{\phi}$. For additional details on the architectures please consider Appendix D.
@@
[…]
@@
where $L$ is a loss function, here we are using the mean-squared error and the cross-entropy loss. To derive our theory, we assume the infinite data limit which allows us to consider a simpler multi-task learning problem with $\mathcal{D}_{\bm{z}}^{\text{query}}=\mathcal{D}_{\bm{z}}^{\text{support}}$ resulting in a single objective
@@
We optimize it in practice by running gradient descent on the task-specific variable $\phi_{{\bm{z}}}$ until convergence for each gradient update of the task-shared parameters $\theta$. In contrast to the bilevel optimization, no second order gradients need to be computed. At inference, we consider a new distribution $\mathcal{P}_{\bm{z}}^{\text{test}}$, for which the support can be disjoint with that of $\mathcal{P}_{\bm{z}}$, and report the expected loss $\mathbb{E}_{{\bm{z}}\sim\mathcal{P}_{\bm{z}}^{\text{test}}}\left[\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})\right]$.
@@ 3 Theory
    No over-parameterization of the student wrt. the teacher. For instance, if the student has more modules available than there are tasks in the training distribution, it could use one module per task, fitting all of them perfectly but remaining ignorant of their shared structure.
@@ 3.1 Multi-task teacher-student setup
In a standard teacher-student setting, a student network is trained to match the output of a teacher given inputs ${\bm{x}}\in\mathbb{R}^{n}$ drawn from some distribution $\mathcal{P}_{\bm{x}}$. Here, we consider a multi-task counterpart to this setting. In addition to the input, we sample a hidden task latent variable $\bm{z}$} from some distribution $\mathcal{P}_{\bm{z}}$ independently from $\bm{x}$}. In the following, we denote all quantities relating to the student with a hat, e.g. $\hat{\Theta}$, and those relating to the teacher without.
@@
where $\psi$ is an activation function, $\bm{a}$} is the task-shared readout vector, and $\Theta=(\Theta^{(m)})_{1\leq m\leq M}$ are the parameter modules parameterizing the hypernetwork. The first layer weight $\bm{W}$} is generated by linearly combining the $(\Theta^{(m)})_{1\leq m\leq M}$ with the task latent variable ${\bm{z}}=(z^{(1)}..z^{(M)})^{\top}$, i.e. ${\bm{W}}(\Theta,{\bm{z}})=\sum_{m}^{M}z^{(m)}\Theta^{(m)}$. The teacher then produces task-conditioned demonstrations as in Equation 3. The task thus has a compositional structure as defined in 2.
[…]
$L(\phi_{{\bm{z}}},\theta;\mathcal{D}_{\bm{z}})=\frac{1}{2}\mathbb{E}_{\bm{x}}\big[\|f({\bm{x}};\hat{{\bm{a}}},\hat{\Theta},\hat{{\bm{z}}})-f({\bm{x}};{\bm{a}},\Theta,{\bm{z}})\|^{2}\big]$ (4)
Figure 2: A Visualization of connected and disconnected task distributions in a teacher with $M=6$, $K=2$. B Module alignment between the student and teacher is high for both continuous and discrete task distributions with connected but not disconnected support. C Module alignment is sensitive to over-parameterization. Numbers denote the factor by which the student dimension is larger than the teacher. Error bars in B,C denote the standard error of the mean over 3 seeds.
@@ 3.2 Identification & Compositional generalization
Let us first study a motivating example for a teacher with $M=6$ modules illustrated in Figure 2A to build some intuition. For that purpose we consider task families associated with binary masks that indicate which sparse subset of $K=2$ modules is present. For instance, the binary mask ${\bm{b}}_{1}=(1,1,0,0,0,0)^{\top}$ indicates that $Z_{1}$ is a task family consisting of latent vectors for which exactly the first and second entry are non-zero, e.g. ${\bm{z}}=(0.6,0.7,0,0,0,0)^{\top}$.
We now consider the case where during training the student observes tasks from four different task families specified by the binary masks ${\bm{b}}_{1}=(1,1,0,0,0,0)^{\top},{\bm{b}}_{2}=(0,1,1,0,0,0)^{\top},{\bm{b}}_{3}=(0,0,0,1,1,0)^{\top},{\bm{b}}_{4}=(0,0,0,0,1,1)^{\top}$. For each task, we assume that the student sees an infinite number of samples covering the whole input space and learns to perfectly fit the teacher. After training we ask the student to solve a new task specified by the task latent code ${\bm{z}}_{\text{?}}=(1,0,0,0,0,1)^{\top}$. Given that it has seen all modules present in this new task, we might expect it to be able to solve it. It turns out that this is not the case since the neuron permutations between modules indexed by $\{1,2,3\}$ and $\{4,5,6\}$ are likely inconsistent as it has never encountered a task that connects these two subsets of modules (also see Appendix A.2.2 for a more detailed explanation). We can avoid this pathology by adding a fifth task family to our set of training tasks given by the binary mask ${\bm{b}}_{5}=(0,0,1,1,0,0)^{\top}$. As we will show now, this leads to compositional generalization and the student will perfectly solve any task including $\bm{z}$$_{\text{?}}$. In fact, the student modules will be a linear combination of the teacher modules which we refer to as linear identification.
@@
$\mathcal{P}_{\bm{z}}$ has compositional support if all dimensions of $\mathbb{R}^{M}$ are covered by at least one of the binary masks ${\bm{b}}_{k}$, i.e. $\sum_{k}^{|Z|}{\bm{b}}_{k}$ has no entry that is 0.
@@
With these two definitions we can now state a simplified version of our main result showing that a perfect student will generalize compositionally. Formal versions of the following theorems as well as their proofs can be found in Appendix A.2.3 and A.2.1, shown for both ReLU as well as a general class of smooth activation functions.
@@
Assuming that $\mathcal{P}_{\bm{z}}$ has compositional and connected support, $\mathcal{P}_{\bm{x}}$ has full support in the input space and the student dimensions match that of the teacher, i.e. $\hat{M}=M\leq n,\hat{h}=h$, then under an additional smoothness condition on $\mathcal{P}_{\bm{z}}$ and non-degeneracy of $(\Theta,{\bm{a}})$, we have that
$\mathbb{E}_{{\bm{z}}\sim\mathcal{P}_{\bm{z}}}\left[\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})\right]=0\quad\implies\quad\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})=0\quad\forall{\bm{z}}\in{\mathbb{R}}^{M},$
i.e. if the student optimizes $\theta$ to fit the teacher on $\mathcal{P}_{\bm{z}}$, it achieves compositional generalization.
The assumptions of compositional, connected support and $h=\hat{h}$ are not only sufficient, but also necessary conditions for the implication to hold (all other things being equal), since we can construct counterexamples, detailed in Appendix A.2.2, when one of the conditions is violated. Furthermore, achieving compositional generalization under the above assumptions is equivalent to linear identification of the teacher modules in the student. For any hidden neuron index $i$, we denote by $\Theta_{i}$ the slice of $\Theta$ of dimension $n\times M$ which maps the task latent variable $\bm{z}$} to ${\bm{w}}_{i}$, i.e. ${\bm{w}}_{i}=\Theta_{i}{\bm{z}}$.
@@
Assuming $\mathcal{P}_{\bm{x}}$ has full support in the input space, $\hat{M}=M\leq n$, $\hat{h}=h$, non-degeneracy of $\Theta$ and $\bm{a}$}, then $\min_{\phi_{\bm{z}}}L(\phi_{\bm{z}},\hat{\Theta};\mathcal{D}_{\bm{z}})=0\quad\forall{\bm{z}}$ is equivalent to the existence of an invertible matrix $\bm{F}$}, a permutation $\sigma$ and a sign flip $\epsilon\in\{-1,1\}^{h}$ such that
$\Theta_{\sigma(i)}=\epsilon_{i}\hat{\Theta}_{i}{\bm{F}}\quad\forall i\in[1,h]$
@@
The linear identification in fact holds in the general setting of $\hat{h}\geq h$, c.f. Appendix A.2.1. So far we have allowed for continuous task families that contain linear combinations of their constituent modules. However, in many interesting settings modules are discrete meaning the task latent variables can only take a finite set of values. Appendix A.2.3 shows that we can capture this case since if the student achieves zero loss on a finite number, $N=\dim(\operatorname{span}(Z_{k}))+1$, of i.i.d. samples of tasks for each $Z_{k}$ from $\mathcal{P}_{\bm{z}}$ it will generalize compositionally with a probability of one.

### v2

@@ Abstract
Many complex tasks can be decomposed into simpler, independent parts. Discovering such underlying compositional structure has the potential to enable compositional generalization. Despite progress, our most powerful systems struggle to compose flexibly. It therefore seems natural to make models more modular to help capture the compositional nature of many tasks. However, it is unclear under which circumstances modular systems can discover hidden compositional structure. To shed light on this question, we study a teacher-student setting with a modular teacher where we have full control over the composition of ground truth modules. This allows us to relate the problem of compositional generalization to that of identification of the underlying modules. In particular we study modularity in hypernetworks representing a general class of multiplicative interactions. We show theoretically that identification up to linear transformation purely from demonstrations is possible without having to learn an exponential number of module combinations. We further demonstrate empirically that under the theoretically identified conditions, meta-learning from finite data can discover modular policies that generalize compositionally in a number of complex environments.  [^1] Code available at https://github.com/smonsays/modular-hyperteacher
@@ 1 Introduction
While monolithic architectures, in which the whole network is engaged for all tasks, are powerful, they lack a built-in mechanism to decompose learned knowledge into modules. This puts them at a potential disadvantage as they might have to learn each of the exponentially many combinations explicitly even if their constituent modules were all previously encountered. For instance, it is unclear to what extent large language models can compositionally generalize beyond the training distribution (Srivastava et al., 2023; Press et al., 2023; Dziri et al., 2023).
This raises the question whether an architecture built from independent modules that can be flexibly combined can more easily extract modular structure from a family of tasks with underlying compositional structure. Typically modular architectures are expressive enough to collapse to the monolithic solution of modeling all task combinations encountered during training and often do so in practice (Shazeer et al., 2017; Kirsch et al., 2018; Rosenbaum et al., 2018; Mittal et al., 2022). Identifying the correct modules is further complicated by the fact that the learning system only ever encounters data generated from mixtures of modules without access to the ground truth decomposition.
Here we aim to illuminate the question what properties of the data generating process allow to discover modular solutions enabling compositional generalization. We do so within the framework of multi-task learning where a learner encounters input/output pairs on a set of tasks composed of few underlying modules. As modular architecture, we consider hypernetworks (Ha et al., 2017), a general form of multiplicative interaction (Jayakumar et al., 2020) which allow to express compositions in weight space used by many multi-task models (Ruvolo & Eaton, 2013; Perez et al., 2018; Dimitriadis et al., 2023; Rame et al., 2023). Our main contributions are as follows:
@@
  • In an extended teacher-student setting, we demonstrate empirically how meta-learning from finite data of sparse module combinations can discover modular solutions that generalize compositionally under the theoretically identified conditions in modular but not monolithic architectures.
  • We show that our results translate to complex settings of learning action-value functions for compositional preferences and learning of behavior policies for compositional goals.
Figure 1: A Diagram contrasting a hypernetwork as a modular architecture to a linear combination of features as a monolithic architecture. B Toy example illustrating the modular teacher-student setting where tasks are drawn from parameter module compositions of a teacher hypernetwork.
@@ 2 Methods
To endow our data generating distributions with compositional structure, for each task we recombine sparse subsets of up to $K$ components from a pool of $M$ modules. Each task can be fully described by a latent code ${\bm{z}}\in\mathbb{R}^{M}$ with $\|{\bm{z}}\|_{0}\leq K$, that specifies the combination of module parameters $(\Theta^{(m)})_{1\leq m\leq M}$ which parameterizes a task-shared function $g$ with parameters $\bm{\omega}({\bm{z}})=\sum^{M}_{m=1}z^{(m)}\Theta^{(m)}$. Being exposed only to demonstrations of such tasks with their underlying modular structure hidden, the naive strategy for a learner would be to learn every combination separately. As the number of modules grows, the quantity of possible combinations increases exponentially, requiring more and more capacity. In contrast, identifying underlying compositional structure not only affords a more compact representation, it also potentially allows the learner to generalize to unseen compositions.
@@
For a learning system to be able to leverage the compositional structure of a family of tasks, its model architecture needs to be able to efficiently represent the corresponding class of compositions. Here we contrast modular and monolithic (non-modular) architectures (see Figure 1A) and investigate under which circumstances they admit functionally modular solutions that reflect the compositional structure of the data generating distribution. To test our models for compositional generalization, we hold-out a number of module combinations when we sample tasks for training. At evaluation time we can then test the ability of all models to solve the held-out out-of-distribution (OOD) tasks.
As our modular architecture of choice we consider hypernetworks (Ha et al., 2017) that allow to express a wide class of multiplicative interactions (Jayakumar et al., 2020). A hypernetwork $h$ is a neural network that generates the parameters $\bm{\omega}$ for a target neural network $g({\bm{x}};\bm{\omega})$ which itself processes the task data. The hypernetwork takes as input a task-specific embedding vector $\bm{\phi}$ and is parameterized by the task-shared parameters $\bm{\theta}$, i.e. $\bm{\omega}=h(\bm{\phi};\bm{\theta})$. We can think of $h$ as a function that maps $\bm{\phi}$ to a selection of flattened module parameters $(\Theta^{(m)})_{1\leq m\leq M}$. In the case of a linear hypernetwork, $h$ is a linear function of $\bm{\phi}$ resulting in a linear combination of the module parameters. In the case of nonlinear hypernetworks, $h$ is a separate neural network whose penultimate layer can be thought of as selecting among the module parameters.
As an example for a monolithic architecture that only supports a less expressive form of composition, we consider the commonly used motif of combining a nonlinear feature encoder shared across tasks with a task-specific linear decoder (e.g. Lee et al., 2019; Bertinetto et al., 2019; Raghu et al., 2020). Here we focus on ANIL (Raghu et al., 2020), i.e. $f({\bm{x}},\bm{\phi},\bm{\theta})=\bm{\phi}^{\top}g({\bm{x}};\bm{\theta})$ where $\bm{\phi}$ is a task-specific readout weight matrix. In addition, we consider the widespread MAML architecture (Finn et al., 2017) with $f({\bm{x}},\bm{\phi},\bm{\theta})=g({\bm{x}};\bm{\phi}(\bm{\theta}))$ where the task-shared parameters $\bm{\theta}$ are used to initialize the task-specific parameters $\bm{\phi}$. For additional details on the architectures please consider Appendix D.
@@
Even if a model supports flexible compositional representations, discovering the hidden latent structure of a task distribution simply from observing demonstrations is difficult. A common paradigm to extract structure shared across tasks is meta-learning which we will consider in the following (Finn et al., 2017; Zhao et al., 2020). Formally, we consider a distribution $\mathcal{P}_{\bm{z}}$ over task latent codes ${\bm{z}}$ which has compositional structure. A given task ${\bm{z}}$ defines a joint distribution over input ${\bm{x}}$ and output ${\bm{y}}$, from which two datasets are sampled, $\mathcal{D}_{\bm{z}}^{\text{support}}=\{x_{i},y_{i}\}_{i=1}^{N^{\text{support}}}$ and $\mathcal{D}_{\bm{z}}^{\text{query}}=\{x^{\prime}_{i},y^{\prime}_{i}\}_{i=1}^{N^{\text{query}}}$. We allow the model to infer the task latent code from the demonstrations in the support set $\mathcal{D}_{\bm{z}}^{\text{support}}$ by optimizing the task-specific parameters $\phi$, obtaining $\phi_{{\bm{z}}}(\theta)$, before querying the model on the query set $\mathcal{D}_{\bm{z}}^{\text{query}}$ to update the task-shared parameters $\theta$. The meta-learning problem can be formalized as a bilevel optimization problem
@@
where $L$ is a loss function; here we are using the mean-squared error and the cross-entropy loss. To derive our theory, we assume the infinite data limit which allows us to consider a simpler multi-task learning problem with $\mathcal{D}_{\bm{z}}:=\mathcal{D}_{\bm{z}}^{\text{query}}=\mathcal{D}_{\bm{z}}^{\text{support}}$ resulting in a single objective
@@
We optimize it in practice by running gradient descent on the task-specific variable $\phi_{{\bm{z}}}$ until convergence for each gradient update of the task-shared parameters $\theta$. In contrast to the bilevel optimization, no second order gradients need to be computed. At inference, we consider a new distribution $\mathcal{P}_{\bm{z}}^{\text{test}}$, for which the support can be disjoint with that of $\mathcal{P}_{\bm{z}}$, and report the expected loss $\mathbb{E}_{{\bm{z}}\sim\mathcal{P}_{\bm{z}}^{\text{test}}}\left[\min_{\phi_{\bm{z}}}\;L(\phi_{\bm{z}},\theta;\mathcal{D}_{\bm{z}})\right]$.
@@ 3 Theory
    No overparameterization of the student wrt. the teacher. For instance, if the student has more modules available than there are tasks in the training distribution, it could use one module per task, fitting all of them perfectly but remaining ignorant of their shared structure.
@@ 3.1 Multi-task teacher-student setup
In a standard teacher-student setting, a student network is trained to match the output of a teacher given inputs ${\bm{x}}\in\mathbb{R}^{n}$ drawn from some distribution $\mathcal{P}_{\bm{x}}$. Here, we consider a multi-task counterpart to this setting (Figure 2A). In addition to the input, we sample a hidden task latent variable ${\bm{z}}$ from some distribution $\mathcal{P}_{\bm{z}}$ independently from ${\bm{x}}$. In the following, we denote all quantities relating to the student with a hat, e.g. $\hat{\Theta}$, and those relating to the teacher without.
@@
where $\psi$ is an activation function, ${\bm{a}}$ is the task-shared readout vector, and $\Theta=(\Theta^{(m)})_{1\leq m\leq M}$ are the parameter modules parameterizing the hypernetwork. The first layer weight ${\bm{W}}$ is generated by linearly combining the $(\Theta^{(m)})_{1\leq m\leq M}$ with the task latent variable ${\bm{z}}=(z^{(1)}..z^{(M)})^{\top}$, i.e. ${\bm{W}}(\Theta,{\bm{z}})=\sum_{m}^{M}z^{(m)}\Theta^{(m)}$. The teacher then produces task-conditioned demonstrations as in Equation 3. The task thus has a compositional structure as defined in Section 2.
The student is also implemented as a linear hypernetwork with the hidden dimension potentially different from the teacher. Crucially, the student only has access to the input ${\bm{x}}$ but does not observe the task latent variable ${\bm{z}}$. We allow the student however to infer a task-specific parameter $\hat{{\bm{z}}}$ of dimension $\hat{M}$ for each task given contextual information which will then condition the forward pass. We estimate $\hat{{\bm{z}}}$ such that it minimizes the objective detailed in Section 2, given demonstrations from the teacher. The student then optimizes the multi-task objective stated in Equation 2, where given the task shared parameters $\theta=(\hat{\Theta},\hat{{\bm{a}}})$, and the task-specific parameters $\phi_{\bm{z}}=\hat{{\bm{z}}}$, the task loss is defined as
$L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})=\frac{1}{2}\mathbb{E}_{\bm{x}}\big{[}\|f({\bm{x}};\hat{{\bm{a}}},\hat{\Theta},\hat{{\bm{z}}})-f({\bm{x}};{\bm{a}},\Theta,{\bm{z}})\|^{2}\big{]}$ (4)
Figure 2: A Schematic of the multi-task teacher-student setup. B Visualization of connected and disconnected task distributions in a teacher with $M=6$, $K=2$. Without the central node support would be disconnected. C Module alignment between the student and teacher is high for both continuous and discrete task distributions with connected but not disconnected support. D Module alignment is sensitive to overparameterization. Numbers denote the factor by which the student dimension is larger than the teacher. Error bars in C,D denote standard error of the mean for 3 seeds.
@@ 3.2 Identification & Compositional generalization
Let us first study a motivating example for a teacher with $M=6$ modules illustrated in Figure 2B to build some intuition. For that purpose we consider task families associated with binary masks that indicate which sparse subset of $K=2$ modules is present. For instance, the binary mask ${\bm{b}}_{1}=(1,1,0,0,0,0)^{\top}$ indicates that $Z_{1}$ is a task family consisting of latent vectors for which exactly the first and second entry are non-zero, e.g. ${\bm{z}}=(0.6,0.7,0,0,0,0)^{\top}$.
We now consider the case where during training the student observes tasks from four different task families specified by the binary masks ${\bm{b}}_{1}=(1,1,0,0,0,0)^{\top},{\bm{b}}_{2}=(0,1,1,0,0,0)^{\top},{\bm{b}}_{3}=(0,0,0,1,1,0)^{\top},{\bm{b}}_{4}=(0,0,0,0,1,1)^{\top}$. For each task, we assume that the student sees an infinite number of samples covering the whole input space and learns to perfectly fit the teacher. After training we ask the student to solve a new task specified by the task latent code ${\bm{z}}_{\text{?}}=(1,0,0,0,0,1)^{\top}$. Given that it has seen all modules present in this new task, we might expect it to be able to solve it. It turns out that this is not the case since the neuron permutations between modules indexed by $\{1,2,3\}$ and $\{4,5,6\}$ are likely inconsistent as it has never encountered a task that connects these two subsets of modules (also see Appendix A.3 for a more detailed explanation). We can avoid this pathology by adding a fifth task family to our set of training tasks given by the binary mask ${\bm{b}}_{5}=(0,0,1,1,0,0)^{\top}$. As we will show now, this leads to compositional generalization and the student will perfectly solve any task including ${\bm{z}}_{\text{?}}$. In fact, the student modules will be a linear combination of the teacher modules which we refer to as linear identification.
@@
$\mathcal{P}_{\bm{z}}$ has compositional support if all dimensions of $\mathbb{R}^{M}$ are covered by at least one of the binary masks ${\bm{b}}_{k}$, i.e. $\sum_{k}^{|Z|}{\bm{b}}_{k}$ has no entry that is 00.
@@
With these two definitions we can now state a simplified version of our main result showing that a perfect student will generalize compositionally. Formal versions of the following theorems as well as their proofs can be found in Appendix A.4 and A.2, shown for both ReLU as well as a general class of smooth activation functions.
@@
Assuming that $\mathcal{P}_{\bm{z}}$ has compositional and connected support, $\mathcal{P}_{\bm{x}}$ has full support in the input space and the student dimensions match that of the teacher, i.e. $-^\hat{M}=M\leq n,\hat{h}=h$, then under an additional smoothness condition on $\mathcal{P}_{\bm{z}}$ and non-degeneracy of $(\Theta,{\bm{a}})$, we have that
$\mathbb{E}_{{\bm{z}}\sim\mathcal{P}_{\bm{z}}}\left[\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})\right]=0\quad\implies\quad\min_{\hat{{\bm{z}}}}\;L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})=0\quad\forall{\bm{z}}\in{\mathbb{R}}^{M},$
i.e. if the student optimizes $(\hat{\Theta},\hat{{\bm{a}}})$ to fit the teacher on $\mathcal{P}_{\bm{z}}$, it achieves compositional generalization.
Furthermore, if the student achieves zero loss on a finite number, $N=\dim(\operatorname{span}(Z_{k}))+1$, of i.i.d. samples of tasks for each $Z_{k}$ from $\mathcal{P}_{\bm{z}}$ it will generalize compositionally almost surely.
This result is noteworthy since it implies that fitting a number of tasks linear in the number of modules suffices to generalize to all exponentially many module combinations.
The assumptions of compositional, connected support and $h=\hat{h}$ are not only sufficient, but also necessary conditions for the implication to hold (all other things being equal), since we can construct counterexamples, detailed in Appendix A.3, when one of the conditions is violated. Furthermore, achieving compositional generalization under the above assumptions is equivalent to linear identification of the teacher modules in the student. For any hidden neuron index $i$, we denote by $\Theta_{i}$ the slice of $\Theta$ of dimension $n\times M$ which maps the task latent variable ${\bm{z}}$ to ${\bm{w}}_{i}$, i.e. ${\bm{w}}_{i}=\Theta_{i}{\bm{z}}$.
@@
Assuming $\mathcal{P}_{\bm{x}}$ has full support in the input space, $\hat{M}=M\leq n$, $-^\hat{h}=h$, non-degeneracy of $\Theta$ and ${\bm{a}}$, then $\min_{\hat{{\bm{z}}}}L(\hat{{\bm{z}}},(\hat{\Theta},\hat{{\bm{a}}});\mathcal{D}_{\bm{z}})=0\quad\forall{\bm{z}}$ is equivalent to the existence of an invertible matrix ${\bm{F}}$, a permutation $\sigma$ and a sign flip $\epsilon\in\{-1,1\}^{h}$ such that
$a_{\sigma{i}}\Theta_{\sigma(i)}=\epsilon_{i}\hat{a}_{i}\hat{\Theta}_{i}{\bm{F}}\quad\forall i\in[1,h]$